The number 2002 is equal to the sum of 7 consecutive natural numbers. What is the
smallest of these 7 numbers?
A) 189 B) 186 C) 283 D) 289 E) 306

Answer by mananth(16949)

let the numbers be x,x+1,x+2,x+3,X=4,x+5,x+6
..
x+x+1+x+2+x+3+X=4+x+5+x+6 =2002
7x+21=2002
7x=2002-21
x=1981/7
x=283
..
The smallest number